# Near-Optimal Regret for Efficient Stochastic Combinatorial Semi-Bandits

### v1

@@ C.1 Comparison with baselines and beyond
CUCB 9.586 9.141 5.743 5.518
EXP3.M 76.213 74.993 43.095 43.169
HYBRID 126.548 133.958 72.772 76.923
Table 2: Semi-bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.5,0.6]$, evaluated on both synthetic and Yelp datasets. Results correspond to subplots (1)(2)(5)(6) in Figure 1.
@@
C.2 Ablation studies and beyond
  • Varying $k$: $k\in\{5,10,15\}$, with $m=30$ fixed.
  • Varying $m$: $m\in\{20,30,40\}$, with $k=10$ fixed.
@@ C.2 Ablation studies and beyond
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $m=20$ $m=30$ $m=40$
@@
CUCB 4765.693 4306.188 3207.724 2156.02 4229.29 6091.72
EXP3.M 4959.489 3525.431 2326.045 1504.01 3437.59 5561.58
HYBRID 6187.287 6140.912 4575.796 3010.17 5349.40 7788.50
[…]
CUCB 4.758 7.531 10.000 7.72 7.62 6.81
EXP3.M 49.400 53.130 52.208 39.50 55.44 64.31
HYBRID 108.027 105.699 101.324 77.68 106.35 132.66
Table 3: Ablation study of semi-bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=10$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 1.
@@
CUCB 5.48 8.08 11.33 8.26 5.93 7.27
EXP3.M 54.72 59.82 64.85 39.58 42.93 58.79
HYBRID 121.78 113.57 114.43 68.54 91.91 111.65
Table 4: Ablation study of semi-bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=10$ fixed) on Yelp dataset.
@@
C.3 Parameter selection for EXP3.M
The performance of EXP3.M is highly sensitive to the choice of its mixing coefficient $\gamma$, which governs the balance between exploration and exploitation. To identify a suitable value for our experimental setting, we conduct an ablation study over three candidate values: $\gamma\in\{0.001,0.01,0.1\}$. Each configuration is executed for a fixed horizon of $T=100000$, repeated over 10 independent runs; results are reported as averages across these repetitions.
As shown in Table 5 and Figure 3, $\gamma=0.01$ consistently achieves the lowest cumulative regret across both initialization settings: $\mathtt{U}(0,0.1)$ and $\mathtt{U}(0.5,0.6)$. Under $\mathtt{U}(0,0.1)$, it reduces regret by up to 82.1% compared to $\gamma=0.001$, and by 60.1% relative to $\gamma=0.1$. Similar patterns are observed under $\mathtt{U}(0.5,0.6)$, where $\gamma=0.01$ also outperforms both alternatives by notable margins. While the runtime remains relatively stable across different $\gamma$ values, their regret outcomes vary significantly. Smaller values like $\gamma=0.001$ lead to insufficient exploration, resulting in poor learning. In contrast, larger values such as $\gamma=0.1$ induce excessive exploration noise, which also degrades performance.
Based on these findings, we adopt $\gamma=0.01$ as the default parameter for all experiments involving EXP3.M, as it strikes the most favorable balance between exploration and exploitation in our stochastic environment.
Metric Algorithm Range $[0,0.1]$ (Synthetic) Range $[0.5,0.6]$ (Synthetic)
$\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$
Regret EXP3.M 19889.88 3569.25 8943.35 18745.12 4223.59 27926.72
Runtime (sec) EXP3.M 48.23 63.86 65.28 53.21 56.35 52.46
Table 5: Ablation study of EXP3.M under different exploration parameters $\gamma$ on the synthetic dataset. Results are shown under two base-arm initial value distributions: (1) $\mathtt{U}(0,0.1)$; (2) $\mathtt{U}(0.5,0.6)$.
C.4 Cascading bandit
Furthermore, we simulate a disjunctive cascading bandit problem with $m=30$ base arms and cardinality $k=10$. In each round, the algorithm recommends a list of $k$ items, and the user examines them sequentially—either in descending or ascending order of attraction probability—clicking on the first attractive item (if any) and ignoring the remaining ones. If none of the items is deemed attractive, the round results in no feedback and is treated as a no-op. The experimental configurations for both datasets (including synthetic and Yelp) and ablation studies in this section follow the same setup as detailed in Section 4 and Appendix C. Particularly, all experiments presented hereafter adopt this disjunctive feedback model. Visual comparisons across all experiments are summarized in Figures 4 (descending) and 5 (ascending).
Descending bandit
[…]
Figure 4: Comparison of CMOSS (blue) with CUCB algorithms under cascading (descending) feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=10$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.5,0.6]$. Shaded regions and error bars denote one standard deviation over 10 runs.
Metric Algorithm Range $[0,0.1]$ Range $[0.5,0.6]$
@@ C.4 Cascading bandit
CUCB 9.313 9.873 4.676 4.560
Table 6: Descending bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.5,0.6]$, evaluated on both synthetic and Yelp datasets. Results correspond to subplots (1)(2)(5)(6) in Figure 4.
@@
Tables 7 and 8 present ablation results on the synthetic and Yelp datasets under the descending-order cascading model, where we vary the cardinality $k$ and the number of base arms $m$. On the synthetic dataset ($\mathtt{U}(0,0.1)$), CMOSS achieves the lowest cumulative regret across all configurations. When varying $k$, CMOSS reduces regret by up to 49.7% compared to CUCB. As $m$ increases, CMOSS continues to outperform CUCB with regret reductions of up to 47.5%. In terms of runtime, CMOSS remains competitive: although slightly slower than CUCB when varying $k$ (e.g., up to 20.7% slower at $k=15$), it is significantly more scalable as $m$ increases. Notably, when $m=40$, CMOSS runs 21.4% faster than CUCB, demonstrating its practical efficiency in larger-scale settings. On the Yelp dataset (rescaled to $[0,0.1]$), similar trends are observed. CMOSS consistently outperforms CUCB in regret across all settings, with improvements of up to 41.7% when varying $k$ and up to 43.9% when varying $m$. In terms of runtime, CMOSS is consistently only marginally slower than CUCB when varying kk—e.g., up to 18.6% slower at k=15k=15—while maintaining stable computational efficiency as $m$ increases. Across all configurations, the runtime gap remains modest: CMOSS is only 20.6% slower when $m=30$, and the gap further narrows to just 14.6% at $m=40$. These results indicate that CMOSS achieves strong regret reduction with only a small and stable runtime overhead, confirming its practical efficiency and scalability. In summary, across all settings, CMOSS maintains reasonable computational cost while scaling better than CUCB, confirming its robustness and real-world applicability.
@@
CUCB 7.571 9.254 11.463 5.257 7.367 6.381
Table 7: Ablation study of descending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=10$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 4.
@@
CUCB 5.842 8.791 12.164 10.435 9.188 10.961
Table 8: Ablation study of descending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=10$ fixed) on Yelp dataset.
$\star$ indicates our proposed method (CMOSS).
Ascending bandit
As shown in Table 9, CMOSS consistently achieves the lowest cumulative regret across both value settings and datasets under the ascending-order cascading model. Under the $[0,0.1]$ initialization, CMOSS reduces regret by 45.7% on the synthetic dataset and 43.8% on the Yelp dataset compared to CUCB. In the $[0.5,0.6]$ initialization, CMOSS maintains lower regret, achieving reductions of 3.1% and 14.8% on the synthetic and Yelp datasets, respectively. In terms of runtime, CMOSS runs slightly slower than CUCB in most cases. Despite this marginal overhead, CMOSS demonstrates a strong trade-off between regret minimization and computational efficiency across both synthetic and real-world environments.
Figure 5: Comparison of CMOSS (blue) with CUCB algorithms under cascading (ascending) feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=10$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.5,0.6]$. Shaded regions and error bars denote one standard deviation over 10 runs.
Metric Algorithm Range $[0,0.1]$ Range $[0.5,0.6]$
Synthetic Yelp Synthetic Yelp
[…]
CUCB 2001.943 2841.852 5.827 8.559
[…]
CUCB 6.688 7.211 3.572 4.239
Table 9: Ascending bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.5,0.6]$, evaluated on both synthetic and Yelp datasets. Results correspond to subplots (1)(2)(5)(6) in Figure 5.
$\star$ indicates our proposed method (CMOSS).
[…]
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $m=20$ $m=30$ $m=40$
[…]
CUCB 3307.608 2272.425 1248.947 1147.413 2077.271 2622.186
[…]
CUCB 4.892 6.920 7.019 6.430 7.986 7.139
Table 10: Ablation study of ascending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=10$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 5.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Yelp) Varying $m$ (Yelp)
$k=5$ $k=10$ $k=15$ $m=20$ $m=30$ $m=40$
[…]
CUCB 3774.767 2706.974 1796.423 1299.960 2550.853 3607.151
[…]
CUCB 4.434 6.303 7.429 6.027 7.616 6.609
Table 11: Ablation study of ascending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=10$ fixed) on Yelp dataset.
$\star$ indicates our proposed method (CMOSS).

### v2

@@ C.4 Cascading bandit
CUCB 12.777 16.246 3.328 8.337
[…]
Table 7: Ablation study of descending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on synthetic dataset within the range $[0,0.1]$. Results correspond to subplots (3)(4)(7)(8) in Figure 4.
@@
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
@@
CUCB 4429.727 3446.924 3184.634 2578.753 1131.007 916.646 1391.695 1952.693 2273.459 2751.108
@@
CUCB 6.574 11.654 16.397 21.287 26.991 24.877 16.767 16.434 14.984 14.915
Table 8: Ablation study of descending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on Yelp dataset within the range $[0,0.1]$.
@@
CUCB 8.215 13.148 16.720 19.717 22.924 25.212 20.810 17.956 17.513 15.967
Ascending bandit
[…]
Figure 5: Comparison of CMOSS (blue) with CUCB algorithms under cascading (ascending) feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=15$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.3,0.4]$.
Table 9: Ascending bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.3,0.4]$, evaluated on both synthetic and Yelp datasets, with fixed $k=10,m=30$. Results correspond to subplots (1)(2)(5)(6) in Figure 5.
@@
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
@@
CUCB 13.395 18.039 10.724 11.739
[…]
Table 10: Ablation study of ascending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 5.
@@
CUCB 6.574 11.654 16.397 21.287 26.991 25.692 19.102 16.068 16.542 16.562
Table 11: Ablation study of ascending bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on Yelp dataset.
@@
CUCB 11.738 13.694 21.308 29.020 31.540 26.714 21.569 19.727 18.980 14.424